# In-and-Out: Algorithmic Diffusion for Sampling Convex Bodies

### v3

@@ 1 Introduction
Samplers for convex bodies are based on Markov chains (see §1.4 for a summary). Their analysis is based on bounding the conductance of the associated Markov chain, which in turn bounds the mixing rate. Analyzing the conductance requires combining delicate geometric arguments with (Cheeger) isoperimetric inequalities for convex bodies. An archetypal example of the latter is the following: for any measurable partition $S_{1},S_{2},S_{3}$ of a convex body $\mathcal{K}\subset\mathbb{R}^{d}$, we have
@@ 1.1 Diffusion for uniform sampling
Under this notation, Algorithm 1 corresponds to a Gibbs sampling scheme from the two marginals of $\pi(x,y)$. To be precise, Line 2 and Line 3 correspond to sampling
@@
We implement the latter step through rejection sampling; if the number of trials in Line 3 hits the threshold $N$, then we halt and declare failure of the algorithm. It is well known that such a Gibbs sampling procedure will ensure the desired stationarity of $\pi(x,y)$. Note that, conditioned on the event that the algorithm does not fail, the resulting iterate will be an unbiased sample from the correct distribution.
@@ 1.2 Results
For any given $\eta,\varepsilon\in(0,1)$, $q\geq 1$, and any convex body $\mathcal{K}$ given by a well-defined membership oracle, there exist choices of parameters $h,N$ such that $\mathsf{In\text{-}and\text{-}Out}$, starting from an $M$-warm distribution, with probability at least $1-\eta$, returns $X\sim\mu$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi^{\mathcal{K}})\leq\varepsilon$. The number of proper steps is $\widetilde{O}(qd^{2}\Lambda\log^{2}\nicefrac{{M}}{{\eta\varepsilon}})$, and the expected total number of membership queries is $\widetilde{O}(qMd^{2}\Lambda\log^{6}\nicefrac{{1}}{{\eta\varepsilon}})$, where $\Lambda$ is the largest eigenvalue of the covariance of $\pi^{\mathcal{K}}$.
Even though our guarantee is in the much stronger “metric” of $\mathcal{R}_{q}$ compared to the $\mathsf{TV}$ guarantees of the $\mathsf{Ball\ walk}$, we do not incur any additional asymptotic complexity. To obtain this result, we choose the following values for the parameters: $h^{-1}=\widetilde{\Theta}(d^{2}\log\frac{qM\Lambda}{\eta}\log\log\frac{1}{\varepsilon})$, $N=\widetilde{\Theta}(\eta^{-1}qMd^{2}\Lambda\log^{5}(1/\varepsilon))$. See Lemma 13 for more details.
Remark 2 (Warm-start generation).
@@
Assume that $\pi^{\mathcal{K}}$ is near-isotropic, i.e., the operator norm of its covariance is $O(1)$. Under the same setting as above, $\mathsf{In\text{-}and\text{-}Out}$ succeeds with probability $1-\eta$, returning $X\sim\mu$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi^{\mathcal{K}})\leq\varepsilon$. The number of proper steps is $\widetilde{O}(qd^{2}\log^{2}\frac{M}{\eta\varepsilon})$, and the expected total number of membership queries is $\widetilde{O}(qMd^{2}\log^{6}\tfrac{1}{\eta\varepsilon})$.
Our analysis will in fact show that the bound on the number of proper steps holds for general non-convex bodies and any feasible start in $\mathcal{K}$. This is deduced under an $M$-warm start in Corollaries 24 and 25. We remark that such a bound for non-convex uniform sampling is not known for the $\mathsf{Ball\ walk}$ or the $\mathsf{Speedy\ walk}$.
Theorem 5.
We have two different convergence results above under (8) and (7). Under (8) we have a doubly-logarithmic dependence on the warmness parameter $M$. On the other hand, using (7), which is weaker than (8) (in general, $C_{\mathsf{PI}}\leq C_{\mathsf{LSI}}$), the dependence on $M$ is logarithmic. We discuss implications of our results further in §1.3.
@@ 1.3 Discussion
Subsequent work.  The ideas and methods of this paper have led to subsequent progress. First, [KZ25] shows that the output guarantee can be strengthened to $\mathcal{R}_{\infty}$ under LSI. Second, the diffusion-based approach is extended to the general logconcave sampling in [KV25b], beyond the uniform distribution. Lastly, the $\mathcal{R}_{\infty}$-warmness requirement has been relaxed to significantly weaker $\mathcal{R}_{\widetilde{O}(1)}$-warmness in [KV25a].
@@ 1.4 Related work
$\mathsf{Hit\text{-}and\text{-}Run}$ is another zeroth-order algorithm that needs no step size; it picks a uniform random line $\ell$ passing a current point, and move to a uniform random point on $\ell\cap\mathcal{K}$. [LV06] shows that, if we define the second moment as $R^{2}:=\mathbb{E}_{\pi^{\mathcal{K}}}[\lvert X-\mathbb{E}X\rvert^{2}]$, then $\mathsf{Hit\text{-}and\text{-}Run}$ requires $O(d^{2}R^{2}\log\frac{M}{\varepsilon})$ queries. Notably, this algorithm has a poly-logarithmic dependence on $M$ as opposed to the $\mathsf{Ball\ walk}$.
@@ 3.2 Contraction and mixing
Proposition 1 (General de Bruijn identity).
@@
We can now deduce contraction under the heat flow.
@@
Let $\mu,\nu$ be probability measures on $\mathbb{R}^{d}$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\nu)<\infty$. Then, for any $t>0$ and $q>1$,
@@
Note that one iteration of $\mathsf{In\text{-}and\text{-}Out}$ corresponds to the composition of two transition kernels $P_{h}$ and $Q_{h}$, where $P_{h}(x,\cdot)=\mathcal{N}(x,hI_{d})$. The forward step ($\mathsf{FH}$) (i.e., transition kernel $P_{h}$) convolves the law $\mu_{k}^{X}$ of the $k$-th iterate with Gaussian $\gamma_{h}$, so the forward step invokes the contraction with rate dependent on $C_{\mathsf{PI}}(\pi^{X})$ or $C_{\mathsf{LSI}}(\pi^{X})$ by Lemma 19. Using the DPI (Lemma 9) to address the effect of $Q_{h}$, we obtain that
$\mathcal{R}_{q}(\mu_{k+1}^{X}\mathbin{\|}\pi^{X})=\mathcal{R}_{q}(\mu_{k+1}^{Y}Q_{h}\mathbin{\|}\pi^{Y}Q_{h})\underset{\text{DPI}}{\leq}\mathcal{R}_{q}(\mu_{k+1}^{Y}\mathbin{\|}\pi^{Y})\leq\frac{\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{LSI}}(\pi^{X}))^{1/q}}\,.$
@@ 3.4 Putting it together
For any given $\eta,\varepsilon\in(0,1)$, $q>1$, $m\in\mathbb{N}$ defined below and any convex body $\mathcal{K}$ given by a well-defined membership oracle, consider $\mathsf{In\text{-}and\text{-}Out}$ (Algorithm 1) with an $M$-warm initial distribution $\mu_{0}^{X}$, step size $h=(2d^{2}\log\frac{9mM}{\eta})^{-1}$, and threshold $N=\widetilde{O}(\frac{mM}{\eta})$. For $\pi^{X}$ the uniform distribution over $\mathcal{K}$,
  • It achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{O}(qd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{2}\frac{M}{\eta\varepsilon})$ iterations. With probability $1-\eta$, the algorithm iterates this many times without failure, using $\widetilde{O}(qMd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{6}\frac{1}{\eta\varepsilon})$ expected number of membership queries in total.
  • For isotropic $\pi^{X}$, with probability $1-\eta$, the algorithm achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ with $m=\widetilde{O}(qd^{2}\log^{2}\frac{M}{\eta\varepsilon})$ iterations, using $\widetilde{O}(qMd^{2}\log^{6}\frac{1}{\eta\varepsilon})$ membership queries in expectation.
@@
We just put together Lemma 13 and Theorem 12. For target accuracy $\varepsilon>0$, we use the $\mathcal{R}_{q}$-decay under (7) for $q\geq 2$ in Theorem 12. The $M$-warm start assumption guarantees $\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})\lesssim\log M$. Due to $C_{\mathsf{PI}}(\pi^{X})=O(\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\,\log d)$ (Lemma 17), $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ if it iterates at least
@@
For isotropic $\pi^{X}$, we have $\operatorname{cov}\pi^{X}=I_{d}$, so the claim immediately follows from $C_{\mathsf{PI}}(\pi^{X})=O(\log d)$. ∎
We now show that the number of proper steps is bounded as claimed for general non-convex bodies and any feasible start in $\mathcal{K}$. We first establish this result under an $M$-warm start (Theorem 5).
Proof of Theorem 5.

### v4

@@ 1 Introduction
Samplers for convex bodies are based on Markov chains (see Sec. §1.4 for a summary). Their analysis is based on bounding the conductance of the associated Markov chain, which in turn bounds the mixing rate. Analyzing the conductance requires combining delicate geometric arguments with (Cheeger) isoperimetric inequalities for convex bodies. An archetypal example of the latter is the following: for any measurable partition $S_{1},S_{2},S_{3}$ of a convex body $\mathcal{K}\subset\mathbb{R}^{d}$, we have
@@ 1.1 Diffusion for uniform sampling
Under this notation, Algorithm 1 without the failure condition corresponds to a Gibbs sampling scheme from the two marginals of $\pi(x,y)$. To be precise, Line 2 and Line 3 correspond to sampling
@@
It is well known that such a Gibbs sampling procedure will ensure the desired stationarity of $\pi(x,y)$. We implement the latter step through rejection sampling; if the number of trials in Line 3 hits the threshold $N$, then we halt and declare failure of the algorithm. Note that the accepted point in Line 3, conditional on success (i.e., number of attempts is at most the cap), is still exactly $\pi^{X|Y=y_{i+1}}$. However, conditioning on “no failure over all $T$ iterations” (or equivalently “re-run-until-success”) re-weights trajectories, so the final output law of the capped INO is not exactly the same as the INO without the cap; see §D for details. Nonetheless, it is $\eta$-close to the intended target $\pi^{X}$ pointwise when the overall failure probability is $\eta$.
Practical implementation via restart.  In practice, one would restart the algorithm if it declares failure before completing the target number of iterations. This yields an always-terminating implementation whose output law equals that of Algorithm 1 conditioned on success. The expected number of restarts is at most $(1-\eta)^{-1}$, so the expected total query complexity increases by a multiplicative factor of only $(1-\eta)^{-1}\leq 2$ for $\eta\leq 1/2$.
@@ 1.2 Results
For any given $\eta,\varepsilon\in(0,1/2)$, $q\geq 2$, and any convex body $\mathcal{K}$ given by a well-defined membership oracle, there exist choices of parameters $h,N$ such that $\mathsf{In\text{-}and\text{-}Out}$, starting from an $M$-warm distribution, with probability at least $1-\eta$, returns $X\sim\mu$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi^{\mathcal{K}})\leq\varepsilon+4\eta$. The number of proper steps is $\widetilde{O}(qd^{2}\Lambda\log^{2}\nicefrac{{M}}{{\eta\varepsilon}})$, and the expected total number of membership queries is $\widetilde{O}(qMd^{2}\Lambda\log^{6}\nicefrac{{1}}{{\eta\varepsilon}})$, where $\Lambda$ is the largest eigenvalue of the covariance of $\pi^{\mathcal{K}}$.
Even though our guarantee is in the much stronger “metric” of $\mathcal{R}_{q}$ compared to the $\mathsf{TV}$ guarantees of the $\mathsf{Ball\ walk}$, we do not incur any additional asymptotic complexity. To obtain this result, we choose the following values for the parameters: $h^{-1}=\widetilde{\Theta}(d^{2}\log\frac{qM\Lambda}{\eta}\log\log\frac{1}{\varepsilon})$, $N=\widetilde{\Theta}(\eta^{-1}qMd^{2}\Lambda\log^{5}(1/\varepsilon))$. See Lemma 15 for more details. The complexity of the restart-on-failure algorithm immediately follows from this result as a simple corollary.
Corollary 4 (Restart-on-failure).
For $\varepsilon\in(0,\frac{1}{10})$ and $q\geq 2$, consider the restart variant that repeats Algorithm 1 from scratch until success with the same parameter choices. Then, the output sample $X\sim\mu^{\prime}$ satisfies $\mathcal{R}_{q}(\mu^{\prime}\,\|\,\pi^{\mathcal{K}})\leq\varepsilon$, and the expected total number of membership queries is $\widetilde{O}(qMd^{2}\Lambda\log^{6}\nicefrac{{1}}{{\varepsilon}})$.
Remark 2 ((Warm-start generation)).
@@
Assume that $\pi^{\mathcal{K}}$ is near-isotropic, i.e., the operator norm of its covariance is $O(1)$. Under the same setting as above, $\mathsf{In\text{-}and\text{-}Out}$ succeeds with probability $1-\eta$, returning $X\sim\mu$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi^{\mathcal{K}})\leq\varepsilon+4\eta$. The number of proper steps is $\widetilde{O}(qd^{2}\log^{2}\frac{M}{\eta\varepsilon})$, and the expected total number of membership queries is $\widetilde{O}(qMd^{2}\log^{6}\tfrac{1}{\eta\varepsilon})$.
Building on the contraction results in [CCSW22, KO25], we show that the bound on the number of proper steps holds for general non-convex bodies and from any feasible start in $\mathcal{K}$. We remark that this bound for non-convex uniform sampling is not known for the $\mathsf{Ball\ walk}$ or the $\mathsf{Speedy\ walk}$.
Corollary 6.
Corollary 7.
We note that, with the step size $h\asymp d^{-2}$, Corollary 7 immediately recovers mixing guarantees matching the ones for the $\mathsf{Speedy\ walk}$ obtained via blocking conductance and the log-Cheeger inequality [KLM06] (see Theorem 28).
@@ 1.3 Discussion
Subsequent work.  The ideas and methods of this paper have led to subsequent progress. First, [KZ25] shows that the output guarantee can be strengthened to $\mathcal{R}_{\infty}$ under LSI, using $\widetilde{O}(d^{2}R^{2}\operatorname{polylog}\frac{1}{\varepsilon\eta})$ queries. Second, the diffusion-based approach is extended to general logconcave sampling in [KV25b], with the query complexity of $\widetilde{O}(d^{2}\,(R^{2}\vee 1)\operatorname{polylog}\frac{1}{\varepsilon\eta})$ for a $\mathcal{R}_{\infty}$ guarantee, beyond the uniform distribution. Lastly, the $\mathcal{R}_{\infty}$-warmness requirement has been relaxed to significantly weaker $\mathcal{R}_{\widetilde{O}(1)}$-warmness in [KV25a], improving the query complexity of warm-start generation from $\widetilde{O}(d^{2}R^{2})$ to $\widetilde{O}(d^{2}R^{3/2}\lVert\operatorname{cov}\pi\rVert^{1/4})$.
@@ 1.4 Related work
$\mathsf{Hit\text{-}and\text{-}Run}$ is another zeroth-order algorithm that needs no step size; it picks a uniform random line $\ell$ passing a current point, and move to a uniform random point on $\ell\cap\mathcal{K}$. [LV06] shows that, if we define the second moment as $R^{2}:=\mathbb{E}_{\pi^{\mathcal{K}}}[\lvert X-\mathbb{E}X\rvert^{2}]$, then $\mathsf{Hit\text{-}and\text{-}Run}$ requires $O(d^{2}R^{2}\log\frac{M}{\varepsilon})$ iterations. Notably, this algorithm has a poly-logarithmic dependence on $M$ as opposed to the $\mathsf{Ball\ walk}$. Chen and Eldan [CE25] further showed that $\mathsf{Hit\text{-}and\text{-}Run}$ mixes in $\widetilde{O}(M^{11}d^{2}/\varepsilon^{11})$ times for isotropic convex bodies.
@@ 3 Analysis
The bias of conditioning on success can be bounded as follows.
Lemma 16 (Bias from no failure).
Let $\mathsf{Succ}$ be the event that the failure is not declared over the required number $m$ iterations, and suppose $\mathbb{P}(\mathsf{Succ})\geq 1-\eta$. Let $\mu_{m}$ be the law of the uncapped $\mathsf{In\text{-}and\text{-}Out}$ output after $m$ iterations, and $\mu_{m}^{\textup{cap}}$ be the law of Algorithm 1 conditioned on $\mathsf{Succ}$. Then, we have:
$\frac{\mathrm{d}\mu_{m}^{\textup{cap}}}{\mathrm{d}\mu_{m}}$ $\leq\frac{1}{1-\eta}\quad\textup{so}\quad\mathcal{R}_{\infty}(\mu_{m}^{\textup{cap}}\|\mu_{m})\leq\log\frac{1}{1-\eta}\,$,
$\mathcal{R}_{q}(\mu_{m}^{\textup{cap}}\,\|\,\pi^{X})$ $\leq\mathcal{R}_{q}(\mu_{m}\|\pi)+\frac{q}{q-1}\,\log\frac{1}{1-\eta}\qquad\text{for }q>1\,.$
@@ 3.2 Contraction and mixing
Proposition 1 (General de Bruijn identity, [KO25]).
@@
We record contraction results under the heat flow and functional inequalities in $q$-Rényi divergence, combining [CCSW22] and Proposition 1.
@@
Let $\mu,\nu$ be probability measures on $\mathbb{R}^{d}$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\nu)<\infty$ and $\mathbb{E}_{\nu}[\lvert\cdot\rvert^{4}]<\infty$. Then, for any $t>0$ and $q>1$,
@@
Note that one iteration of $\mathsf{In\text{-}and\text{-}Out}$ corresponds to the forward and backward step. The forward step ($\mathsf{FH}$) (i.e., transition kernel $P_{h}=\mathcal{N}(x,hI_{d})$) convolves the law $\mu_{k}^{X}$ of the $k$-th iterate with Gaussian $\gamma_{h}$, so the forward step invokes the contraction with rate dependent on $C_{\mathsf{PI}}(\pi^{X})$ or $C_{\mathsf{LSI}}(\pi^{X})$ by Lemma 22. Using the DPI (Lemma 11) to address the effect of the backward step ($\mathsf{BH}$), we obtain that
$\mathcal{R}_{q}(\mu_{k+1}^{X}\mathbin{\|}\pi^{X})\underset{\text{DPI}}{\leq}\mathcal{R}_{q}(\mu_{k+1}^{Y}\mathbin{\|}\pi^{Y})\leq\frac{\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{LSI}}(\pi^{X}))^{1/q}}\,.$
@@ 3.3 Failure probability and wasted steps
Lastly, we prove a simple lemma showing that the bias due to conditioning on success is bounded in terms of the failure probability $\eta$.
Proof of Lemma 16.
Let $X$ be the uncapped $\mathsf{In\text{-}and\text{-}Out}$ output after the desired number $m$ of iterations, and $\mathsf{Succ}$ be the event that the failure is not declared until $m$ iterations. Let $\mu_{m}^{X}$ be the law of $X$ and $\mu_{m}^{\textup{cap}}$ be the law of $X$ conditioned on $\mathsf{Succ}$. Then, by Bayes’ rule,
$\frac{\mathrm{d}\mu_{m}^{\textup{cap}}}{\mathrm{d}\mu_{m}}(x)=\frac{\mathbb{P}(X=x\,|\,\mathsf{Succ})}{\mathbb{P}(X=x)}=\frac{\mathbb{P}(\mathsf{Succ}\,|\,X=x)\,\mathbb{P}(X=x)}{\mathbb{P}(\mathsf{Succ})\,\mathbb{P}(X=x)}\leq\frac{1}{\mathbb{P}(\mathsf{Succ})}\leq\frac{1}{1-\eta}\,.$
Lastly, using the weak triangle inequality for Rényi divergence,
$\mathcal{R}_{q}(\mu_{m}^{\textup{cap}}\,\|\,\pi)\leq\frac{q}{q-1}\,\mathcal{R}_{\infty}(\mu_{m}^{\textup{cap}}\,\|\,\mu_{m})+\mathcal{R}_{q}(\mu_{m}\,\|\,\pi)\leq\mathcal{R}_{q}(\mu_{m}\,\|\,\pi)+\frac{q}{q-1}\log\frac{1}{1-\eta}\,$,
which completes the proof. ∎
@@ 3.4 Putting it together
For any given $\eta,\varepsilon\in(0,1/2)$, $q\geq 2$, $m\in\mathbb{N}$ defined below and any convex body $\mathcal{K}$ given by a well-defined membership oracle, consider $\mathsf{In\text{-}and\text{-}Out}$ (Algorithm 1) with an $M$-warm initial distribution $\mu_{0}^{X}$, step size $h=(2d^{2}\log\frac{9mM}{\eta})^{-1}$, and threshold $N=\widetilde{O}(\frac{mM}{\eta})$. For $\pi^{X}$ the uniform distribution over $\mathcal{K}$, let $\mu_{m}^{X}$ be the law of the $m$-th iterate of $\mathsf{In\text{-}and\text{-}Out}$ conditioned on the success.
  • With probability $1-\eta$, the algorithm iterates this many times without failure, using $\widetilde{O}(qMd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{6}\frac{1}{\eta\varepsilon})$ expected number of membership queries in total and achieving $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon+4\eta$ after $m=\widetilde{O}(qd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{2}\frac{M}{\eta\varepsilon})$ iterations.
  • For isotropic $\pi^{X}$, with probability $1-\eta$, the algorithm achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon+4\eta$ with $m=\widetilde{O}(qd^{2}\log^{2}\frac{M}{\eta\varepsilon})$ iterations, using $\widetilde{O}(qMd^{2}\log^{6}\frac{1}{\eta\varepsilon})$ membership queries in expectation.
@@
We just put together Lemma 15 and Theorem 14. For target accuracy $\varepsilon>0$, we use the $\mathcal{R}_{q}$-decay under (9) for $q\geq 2$ in Theorem 14. The $M$-warm start assumption guarantees $\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})\lesssim\log M$. Due to $C_{\mathsf{PI}}(\pi^{X})=O(\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\,\log d)$ (Lemma 20), $\mathsf{In\text{-}and\text{-}Out}$ “without the cap” can achieve $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ if it iterates at least
@@
For isotropic $\pi^{X}$, we have $\operatorname{cov}\pi^{X}=I_{d}$, so the claim immediately follows from $C_{\mathsf{PI}}(\pi^{X})=O(\log d)$. Lastly, as noted in Lemma 16, the conditional-on-success output distribution of $\mathsf{In\text{-}and\text{-}Out}$ introduces a small bias, bounded by $2\eta$. ∎
We show that the restart variant increases the expected total query complexity by at most a factor of $(1-\eta)^{-1}$.
Proof of Corollary 4.
Since the success probability of $m$ iterations without failure is at least $1-\eta$, the expected number of restarts is $(1-\eta)^{-1}$. Setting $4\eta=\varepsilon(<1/10)$ and replacing $\varepsilon$ by $\varepsilon/2$ in Theorem 26, the conclusion follows. ∎
We now bound the number of proper steps for general non-convex bodies and any feasible start in convex $\mathcal{K}$. We first record the following under an $M$-warm start, as a corollary of the contraction results in [CCSW22, KO25].
Proposition 2.
Proof.